OpenStudy (anonymous):

hi i have an urgent math problem, find the domain and range of the function, y=-5/81x^2+20 use x for domain and y for range, thanks

OpenStudy (inkyvoyd):

First all, you mean (-5/81)x^2+20=y?

OpenStudy (anonymous):

yes the -5 is over the 81x^2+20=y

OpenStudy (anonymous):

\(y=\dfrac{-5}{81x^2+20}\)?

OpenStudy (inkyvoyd):

You want the domain and range right? Do you know what a domain and a range is?

OpenStudy (anonymous):

to nbouscal: that is correct but the +20 is not under the vaculum bar